Calculate the number of lines per millimeter required for a first order diffraction line at 500nm to be observed if the reflection angle is 10 deg and the incident angle is 60 deg
Answer #1

[ sin( thetam) + sin (thetai) ] = m * n * lambda

where thetam = reflection angle at mth mode , thetai = incident angle at i th mode , m = mode , n = number of lines per mm on the grating and lambda is wavelength of light on grating

sin(10) + sin(60) = 1 * n * 500 x 10^-6 mm

1.04 = 1 * n * 5 x 10^-4 mm

n = 2.08 x 10^3 lines per mm